OpenStudy (anonymous):

write the slope intercept form of the equation of a line having: m=2 and y-intercept -4

OpenStudy (anonymous):

y=mx+b m=slope b=y-intercept

OpenStudy (anonymous):

Can you figue it out from there?

OpenStudy (anonymous):

just did

OpenStudy (anonymous):

Great! What did you come up with?

OpenStudy (anonymous):

\[y=2x-4\]

OpenStudy (anonymous):

Correct!

OpenStudy (anonymous):

write the slope-intersept form of the equation of the line passing through the point(-5, 1) and perpendicular to the line y=4/3x=4

OpenStudy (anonymous):

srry im in algebra 2 but im doing a review sheet and cant remember stuff

OpenStudy (anonymous):

Oh sorry i can't help you with that one! I have trouble myself with perpendicular lines!

OpenStudy (anonymous):

its fine its just confusing after spending a whole year in geometry

OpenStudy (anonymous):

No problem! Wish I could help you out there!